Which of the following is the probability of rolling an even number on a 6-sided number cube and flipping a coin twice and getting two head in a row?
1/2
1/8
1/4
1/3

Answer by fractalier(6550)

To get an even number on the die is 3/6 (or 1/2) and getting two heads in a row is (1/2)(1/2) = 1/4.
To have both events happen is (1/2)(1/4) = 1/8